5 years ago radha age was 7 times of her daughter.her age will be 3 times of her daughter in 5 years.then what is the present age of radha?
Answer by mananth(16946)

5 years ago radha age was 7 times of her daughter.her age will be 3 times of her daughter in 5 years.then what is the present age of radha?

let daughter's age =x
Radha's age =y
5 years ago
y-5=7(x-5)
y-5=7x-35
y-7x=-30........................(1)
5 years later
y+5=3(x+5)
y+5=3x+15
y-3x=10.........................(2)
Eliminate y
multiply (1)by -1
Multiply (2) by 1
7 x -1 y = 30
-3 x + 1 y = 10
Add the two equations
4 x = 40
/ 4
x = 10
plug value of x in (1)
-7 x + 1 y = -30
-70 + y = -30
y = -30 + 70
y = 40
y = 40
x= 10
y= 40
m.ananth@hotmail.ca